# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ 3.1 Formulation of the model with periodic boundary conditions: statistical structure of Fourier modes
For practical and numerical reasons, we formulate a version of the Gaussian model described in the former section on the three dimensional torus of size $L_{\text{\tiny{tot}}}$: the set $L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}$ with periodic boundary conditions as has been done in DNS. The model is Gaussian, it is thus fully determined in the statistically homogeneous and stationary regime by its covariance function (34). The respective velocity field will thus thus $L_{\text{\tiny{tot}}}$-periodic $u^{\nu}_{i}(t,\mathbf{x})=u^{\nu}_{i}(t,\mathbf{x}+L_{\text{\tiny{tot}}}\hat{e}_{p})$ in any Cartesian direction $p=1$, 22 or $3$. Let us mention that future simulations will be performed using $L_{\text{\tiny{tot}}}=2\pi$ in order to be consistent with the DNS. A fundamental consequence of considering a periodic setting is the possibility to expand the velocity field using a Fourier series, i.e.
$u^{\nu}_{i}(t,\mathbf{x})=\frac{1}{L^{3}_{\text{\tiny{tot}}}}\sum_{\mathbf{n}\in\mathbb{Z}^{3}}e^{2i\pi\mathbf{k}_{\mathbf{n}}\cdot\mathbf{x}}\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}}),$ (47)
with Fourier modes $\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$, depending on $\mathbf{k}_{\mathbf{n}}=\mathbf{n}/L_{\text{\tiny{tot}}}$ for $\mathbf{n}=(n_{1},n_{2},n_{3})\in\mathbb{Z}^{3}$. The Fourier modes are defined by
$\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})=\int_{[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}e^{-2i\pi\mathbf{k}_{\mathbf{n}}\cdot\mathbf{x}}u^{\nu}_{i}(t,\mathbf{x}){\mathrm{d}}\mathbf{x}.$ (48)
It can be checked that the variance of the Fourier modes is finite. This solves the issue exhibited in the former Section, where Fourier modes in the whole space need to be interpreted in a distributional sense. Since the velocity field $\mathbf{u}^{\nu}$ is real-valued, Fourier modes are hermitian symmetric, i.e.
$\overline{\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})}=\widehat{u}^{\nu}_{i}(t,-\mathbf{k}_{\mathbf{n}})=\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{-\mathbf{n}}).$ (49)
This implies that the Fourier mode for $\mathbf{n}=(0,0,0)$ is real-valued.
The purpose of this Section is the formulation in a periodic setting of the statistical quantities defined in the whole space in Section 2.4. To do so, let us begin by defining the spatio-temporal covariance of the velocity field $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ as
$\mathcal{C}_{ij}^{\nu}(\tau,\ell)=\mathbb{E}\left[u_{i}^{\nu}(t,\mathbf{x})u_{j}^{\nu}(t+\tau,\mathbf{x}+\ell)\right].$ (50)
The definition is identical to which is very similar to (34), however $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ given in (50) is periodic. If it is furthermore required that $\mathcal{C}_{ij}^{\nu}$ (50) characterizes a statistically homogeneous and stationary incompressible velocity field, the right-hand side of (50) is thus independent on the position $\mathbf{x}$ and time $t$. The covariance structure of the Fourier modes is then written as
@@
$\mathcal{C}_{ij}^{\nu}(\tau,\ell)=\frac{1}{2L^{3}_{\text{\tiny{tot}}}}\sum_{\mathbf{n}\in\mathbb{Z}^{3}}e^{2i\pi\mathbf{k}_{\mathbf{n}}\cdot\ell}F(\tau/T_{k_{\mathbf{n}}})E_{\nu}^{\text{\tiny{E}}}(k_{\mathbf{n}})\widehat{P}_{ij}(\mathbf{k}_{\mathbf{n}}),$ (53)
which is obtained from (50).
@@
where the dissipative cut-off $\eta_{d}$ would become proportional to the Kolmogorov dissipative length scale $\eta_{K}=O\left((\nu^{3}/\varepsilon)^{1/4}\right)$ only when $H=1/3$ (12). In the following, we will consider the length scale $\eta_{d}$ as a free parameter. As we have seen in Section 2.1, imposing a given PSD in the longitudinal case $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (10) implies for statistical isotropic reasons a form of the PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ in dimension $d=3$ through the relation provided in (9), giving a particular simple form in the asymptotic limit of vanishing viscosity $E^{\text{\tiny{E}}}(k)$ (13). Using any $0<H<1$, instead of the particular value $H=1/3$, the one-dimensional spectrum (54) leads to, through the relation (9),
@@
$S_{2}^{\text{\tiny{long}}}(\ell)$ $=\lim_{L_{\text{\tiny{tot}}}\to\infty}\lim_{\nu\to 0}\mathbb{E}\left[\left(\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}\right)^{2}\right]=2\int_{k\in\mathbb{R}}\left(1-e^{2i\pi k|\ell|}\right)E^{\text{\tiny{E,long}}}(k){\mathrm{d}}k$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}4D_{2}|\ell|^{2H}\int_{k=0}^{\infty}\left(1-\cos(2\pi k)\right)k^{-(2H+1)}{\mathrm{d}}k$
$=2D_{2}(2\pi)^{2H}\cos(\pi H)\frac{\Gamma(2-2H)}{H(1-2H)}|\ell|^{2H}.$ (58)
Using the well-known relation $\Gamma(z+1)=z\Gamma(z)$, one checks that when $H=1/3$, the prediction (3.1) coincides with (2.1). Concerning the increment of a velocity component (23) and the second-order structure function (24), we derive under the assumption of statistical isotropy and incompressibility a generalization of (25): one has
$S_{2,ij}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}S_{2}^{\text{\tiny{long}}}(|\ell|)\left(-H\frac{\ell_{i}\ell_{j}}{|\ell|^{2}}+(1+H)\delta_{ij}\right),$ (59)
@@
$S_{2}(\ell)\equiv S_{2,ii}(\ell)=\lim_{L_{\text{\tiny{tot}}}\to\infty}\lim_{\nu\to 0}\mathbb{E}\left[\left|\delta_{\ell}\mathbf{u}^{\nu}\right|^{2}\right]$ $=2\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}(3+2H)S_{2}^{\text{\tiny{long}}}(|\ell|).$ (60)
Correspondingly, the characteristic time scale $\tau_{\text{c}}(|\ell|)$ (29) reads, using the expression provided in (2.4) with the generalized forms of $E^{\text{\tiny{E}}}(|\mathbf{k}|)$ (55) and $T_{\mathbf{k}}$ (56), and for any appropriate function $F$, in the limit of an infinitely large periodic box allowing to use the correspondance depicted in (57),
$\lim_{L_{\text{\tiny{tot}}}\to\infty}\tau_{\text{c}}(|\ell|)$ $=\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)T_{\mathbf{k}}E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\ell}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (61)
At small scales, the characteristic time scale $\tau_{\text{c}}(|\ell|)$ (61) undergoes a transition depending on the values of $H$ and $\beta$, following the behaviors
$\lim_{L_{\text{\tiny{tot}}}\to\infty}\tau_{\text{c}}(|\ell|)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}\frac{|\ell|^{2\beta}}{D_{3}}\frac{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1-2\beta}{\mathrm{d}}\rho}{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1}{\mathrm{d}}\rho}\int_{0}^{\infty}F(s){\mathrm{d}}s\mbox{ for }H<1-\beta,$ (62)
@@
$\lim_{L_{\text{\tiny{tot}}}\to\infty}$ $\tau_{\text{c}}(|\ell|)$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\ell|\to 0}}\frac{4\pi^{2}}{3}\frac{|\ell|^{2(1-H)}}{D_{3}}\frac{\int_{\rho\geq 0}\rho^{6}\sqrt{\rho^{2}+1/L^{2}}^{-2H-5-2\beta}{\mathrm{d}}\rho}{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1}{\mathrm{d}}\rho}\int_{0}^{\infty}F(s){\mathrm{d}}s\mbox{ for }H>1-\beta.$ (63)
@@
3.2 Causal evolution of the Fourier modes
Let us now discuss the time evolution of the Fourier modes $\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$ (48), such that the correlation structure is given by (51), leading to the expected covariance structure $\mathcal{C}_{ij}^{\nu}(\tau,\ell)$ given in (53). This cannot be done in full generality for any function $F$ since we are asking for a causal evolution. Actually, very few choices for $F$ can be achieved for a given causal evolution, and we will only explore a very limited set of them. As far as we are concerned, we will present choices of $F$ which give a causal evolution.
3.2.1 Fourier modes of a white noise
Before proceeding, we define a Gaussian white noise which is the building block of Gaussian stochastic processes and its Fourier series. A Gaussian white noise, and more precisely a periodic Gaussian white vector over $\mathbf{x}\in[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}$, will be denoted ${\mathrm{d}}W(\mathbf{x})$”, which is put in quotes because it has no meaning locally at a given position $\mathbf{x}$, and needs to be properly defined, to be integrated against a deterministic test function as defined in (3.2.1). Component-wise, a Gaussian white vector ${\mathrm{d}}W_{i}(\mathbf{x})$ for $i=1,2$ or $3$ is a Gaussian random distribution which is characterized by, for arbitrary and appropriate deterministic functions $f$ and $g$, the average
$\mathbb{E}\left[\int_{\mathbf{x}\in[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}f(\mathbf{x}){\mathrm{d}}W_{i}(\mathbf{x})\right]=0$ (64)
and its covariance
$\mathbb{E}\left[\int_{\mathbf{x}\in[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}f(\mathbf{x}){\mathrm{d}}W_{i}(\mathbf{x})\right$. $\left.\int_{\mathbf{y}\in[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}g(\mathbf{y}){\mathrm{d}}W_{j}(\mathbf{y})\right]$
$=\delta_{ij}\int_{\mathbf{x}\in[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}f(\mathbf{x})g(\mathbf{x}){\mathrm{d}}\mathbf{x}.$ (65)
Fourier modes (48) of the vector white noise will be denoted by $\widehat{{\mathrm{d}}W}(\mathbf{k}_{\mathbf{n}})$ and are defined by
$\widehat{{\mathrm{d}}W}_{i}(\mathbf{k}_{\mathbf{n}})=\int_{[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}e^{-2i\pi\mathbf{k}_{\mathbf{n}}\cdot\mathbf{x}}{\mathrm{d}}W_{i}(\mathbf{x}).$ (66)
Interestingly, whereas the white noise in the physical space has a distributional nature, its Fourier modes can be understood in a classical sense. They are complex Gaussian random variables, Hermitian symmetric, fully characterized by their average, which turns out to be zero, and their covariance structure, which reads
$\mathbb{E}\left[\widehat{{\mathrm{d}}W}_{i}(\mathbf{k}_{\mathbf{n}})\widehat{{\mathrm{d}}W}_{j}(\mathbf{k}_{\mathbf{m}})\right]=L^{3}_{\text{\tiny{tot}}}\delta_{i,j}\delta_{\mathbf{n},-\mathbf{m}}^{(3)},$ (67)
and
$\mathbb{E}\left[\widehat{{\mathrm{d}}W}_{i}(\mathbf{k}_{\mathbf{n}})\overline{\widehat{{\mathrm{d}}W}_{j}(\mathbf{k}_{\mathbf{m}})}\right]=L^{3}_{\text{\tiny{tot}}}\delta_{i,j}\delta_{\mathbf{n},\mathbf{m}}^{(3)}.$ (68)
As far as we are concerned, we furthermore have to add a temporal dependence to this white noise and define the space-time vector white noise ${\mathrm{d}}W_{i}(t,\mathbf{x})$”, which furthermore includes independent in time instances, which again have to be considered in a distributional manner, i.e., it has to be integrated against test functions, in a similar fashion as has been done in (64) and (3.2.1). The respective Fourier mode (48) of the vector white noise will be denoted $\widehat{{\mathrm{d}}W}_{i}(t,\mathbf{k}_{\mathbf{n}})$” as
$``\widehat{{\mathrm{d}}W}_{i}(t,\mathbf{k}_{\mathbf{n}})"=\int_{[-L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}}e^{-2i\pi\mathbf{k}_{\mathbf{n}}\cdot\mathbf{x}}{\mathrm{d}}W_{i}(t,\mathbf{x}).$ (69)
Once again, the quotes refer to its distributional nature in time. Because the Fourier mode is a linear operation, $\widehat{{\mathrm{d}}W}_{i}(t,\mathbf{k}_{\mathbf{n}})$” remains a zero-average Gaussian distribution, but it is complex-valued and Hermitian symmetric. For the sake of simplicity, we write in a formal manner

### v2

@@ 3.1 Formulation of the model with periodic boundary conditions: statistical structure of Fourier modes
For practical and numerical reasons, we formulate a version of the Gaussian model described in the former section on the three dimensional periodic space $L_{\text{\tiny{tot}}}/2\ ;\ L_{\text{\tiny{tot}}}/2]^{3}$ of period $L_{\text{\tiny{tot}}}$, as has been done in DNS. The model is Gaussian, it is thus fully determined in the statistically homogeneous and stationary regime by its covariance function (34). The respective velocity field will thus thus $L_{\text{\tiny{tot}}}$-periodic $u^{\nu}_{i}(t,\mathbf{x})=u^{\nu}_{i}(t,\mathbf{x}+L_{\text{\tiny{tot}}}\hat{e}_{p})$ in any Cartesian direction $p=1$, 22 or $3$. Let us mention that future simulations will be performed using $L_{\text{\tiny{tot}}}=2\pi$ in order to be consistent with the DNS. A fundamental consequence of considering a periodic setting is the possibility to expand the velocity field using a Fourier series and the respective Fourier modes $\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$, depending on $\mathbf{k}_{\mathbf{n}}=\mathbf{n}/L_{\text{\tiny{tot}}}$ for $\mathbf{n}=(n_{1},n_{2},n_{3})\in\mathbb{Z}^{3}$. It can then be checked that the variance of the Fourier modes is finite. This solves the issue exhibited in the former Section, where Fourier modes in the whole space need to be interpreted in a distributional sense. Since the velocity field $\mathbf{u}^{\nu}$ is real-valued, Fourier modes are hermitian symmetric, i.e. $\overline{\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})}=\widehat{u}^{\nu}_{i}(t,-\mathbf{k}_{\mathbf{n}})=\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{-\mathbf{n}})$. This implies that the Fourier mode for $\mathbf{n}=(0,0,0)$ is real-valued.
The purpose of this Section is the formulation in a periodic setting of the statistical quantities defined in the whole space in Section 2.4. To do so, let us begin by defining the spatio-temporal covariance of the velocity field $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ as
$\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})=\mathbb{E}\left[u_{i}^{\nu}(t,\mathbf{x})u_{j}^{\nu}(t+\tau,\mathbf{x}+\boldsymbol{\ell})\right].$ (47)
The definition is identical to which is very similar to (34), however $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ given in (47) is periodic. If it is furthermore required that $\mathcal{C}_{ij}^{\nu}$ (47) characterizes a statistically homogeneous and stationary incompressible velocity field, the right-hand side of (47) is thus independent on the position $\mathbf{x}$ and time $t$. The covariance structure of the Fourier modes is then written as
@@
$\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})=\frac{1}{2L^{3}_{\text{\tiny{tot}}}}\sum_{\mathbf{n}\in\mathbb{Z}^{3}}e^{2i\pi\mathbf{k}_{\mathbf{n}}\cdot\boldsymbol{\ell}}F(\tau/T_{k_{\mathbf{n}}})E_{\nu}^{\text{\tiny{E}}}(k_{\mathbf{n}})\widehat{P}_{ij}(\mathbf{k}_{\mathbf{n}}).$ (50)
@@
where the dissipative cut-off $\eta_{d}$ is a parameter of the formulation. Let us recall that the usual phenomenology of turbulence, as depicted in a self-consistent manner in Section 2.1 is recovered only for $H=1/3$. In the following, we will consider the length scale $\eta_{d}$ as a free parameter. As we have seen in Section 2.1, imposing a given PSD in the longitudinal case $E_{\nu}^{\text{\tiny{E,long}}}(k)$ (10) implies for statistical isotropic reasons a form of the PSD $E_{\nu}^{\text{\tiny{E}}}(k)$ in dimension $d=3$ through the relation provided in (9), giving a particular simple form in the asymptotic limit of vanishing viscosity $E^{\text{\tiny{E}}}(k)$ (13). Using any $0<H<1$, instead of the particular value $H=1/3$, the one-dimensional spectrum (51) leads to, through the relation (9),
@@
$S_{2}^{\text{\tiny{long}}}(\ell)$ $=\lim_{L_{\text{\tiny{tot}}}\to\infty}\lim_{\nu\to 0}\mathbb{E}\left[\left(\delta_{\ell}^{\text{\tiny{long}}}\mathbf{u}^{\nu}\right)^{2}\right]=2\int_{k\in\mathbb{R}}\left(1-e^{2i\pi k\ell}\right)E^{\text{\tiny{E,long}}}(k){\mathrm{d}}k$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}2D_{2}(2\pi)^{2H}\cos(\pi H)\frac{\Gamma(2-2H)}{H(1-2H)}\ell^{2H}.$ (55)
Using the well-known relation $\Gamma(z+1)=z\Gamma(z)$, one checks that when $H=1/3$, the prediction (3.1) coincides with (2.1). Concerning the increment of a velocity component (23) and the second-order structure function (24), we derive under the assumption of statistical isotropy and incompressibility a generalization of (25), one has :
$S_{2,ij}(\boldsymbol{\ell})\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}S_{2}^{\text{\tiny{long}}}(\ell)\left(-H\frac{\ell_{i}\ell_{j}}{\ell^{2}}+(1+H)\delta_{ij}\right),$ (56)
@@
$S_{2}(\ell)\equiv S_{2,ii}(\ell)=\lim_{L_{\text{\tiny{tot}}}\to\infty}\lim_{\nu\to 0}\mathbb{E}\left[\left|\delta_{\ell}\mathbf{u}^{\nu}\right|^{2}\right]$ $=2\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}(3+2H)S_{2}^{\text{\tiny{long}}}(\ell).$ (57)
Correspondingly, the characteristic time scale $\tau_{\text{c}}(\ell)$ (29) reads, using the expression provided in (2.4) with the generalized forms of $E^{\text{\tiny{E}}}(|\mathbf{k}|)$ (52) and $T_{\mathbf{k}}$ (53), and for any appropriate function $F$, in the limit of an infinitely large periodic box allowing to use the correspondence depicted in (54),
$\lim_{L_{\text{\tiny{tot}}}\to\infty}\tau_{\text{c}}(\ell)$ $=\frac{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)T_{\mathbf{k}}E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}{\int_{\mathbf{k}\in\mathbb{R}^{3}}\left(1-e^{2i\pi\mathbf{k}\cdot\boldsymbol{\ell}}\right)E^{\text{\tiny{E}}}(|\mathbf{k}|){\mathrm{d}}\mathbf{k}}\int_{0}^{\infty}F(s){\mathrm{d}}s.$ (58)
At small scales, the characteristic time scale $\tau_{\text{c}}(\ell)$ (58) undergoes a transition depending on the values of $H$ and $\beta$, following the behaviors
$\lim_{L_{\text{\tiny{tot}}}\to\infty}\tau_{\text{c}}(\ell)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}\frac{\ell^{2\beta}}{D_{3}}\frac{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1-2\beta}{\mathrm{d}}\rho}{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1}{\mathrm{d}}\rho}\int_{0}^{\infty}F(s){\mathrm{d}}s\mbox{ for }H<1-\beta,$ (59)
@@
$\lim_{L_{\text{\tiny{tot}}}\to\infty}$ $\tau_{\text{c}}(\ell)$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{\ell\to 0}}\frac{4\pi^{2}}{3}\frac{\ell^{2(1-H)}}{D_{3}}\frac{\int_{\rho\geq 0}\rho^{6}\sqrt{\rho^{2}+1/L^{2}}^{-2H-5-2\beta}{\mathrm{d}}\rho}{\int_{\rho\geq 0}\left(1-\frac{\sin(2\pi\rho)}{2\pi\rho}\right)\rho^{-2H-1}{\mathrm{d}}\rho}\int_{0}^{\infty}F(s){\mathrm{d}}s\mbox{ for }H>1-\beta.$ (60)
@@
3.2 Markovian evolution of the Fourier modes
Let us now discuss the time evolution of the Fourier modes $\widehat{u}^{\nu}_{i}(t,\mathbf{k}_{\mathbf{n}})$, such that the correlation structure is given by (48), leading to the expected covariance structure $\mathcal{C}_{ij}^{\nu}(\tau,\boldsymbol{\ell})$ given in (50). This cannot be done in full generality for any function $F$ since we are asking for a Markovian evolution. Actually, very few choices for $F$ can be achieved for a given Markovian evolution, and we will only explore a very limited set of them. As far as we are concerned, we will present choices of $F$ which give a Markovian evolution in which additional underlying layers will be involved.
Before proceeding, we have to define a space-periodic and temporal Gaussian white noise ${\mathrm{d}}W_{i}(t,\mathbf{x})$", a notation that we put in quotes because it has to be considered as a distribution. More precisely, let us focus on its Fourier modes $\widehat{{\mathrm{d}}W}_{i}(t,\mathbf{k}_{\mathbf{n}})$". Being complex Gaussian and Hermitian, it is eventually fully determined by the covariance function